40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3214     Entry time: Wed Jul 14 11:32:36 2010
Author: Sharmila, Katharine 
Type: Update 
Category: WIKI-40M Update 
Subject:  

Weekly update:

We correctly connected our circuit to power source to verify that it was functional and that our LED in the shadow sensor turned on.  It did, but we also noticed a funky signal from the LED driver.  We continued to attempt 1x1 levitation, but determined that the temporary flag we were using out of convenience (a long, thin cylindrical magnet) was weakly attracted to residually magnetized OSEM components.  We then switched to an aluminum screw as our flag.

We resoldered and applied heat shrink to the wires connecting our coil to the BNC terminal, since they were falling apart.

We sat down with Rana and removed circuit components in the LED drive part by part to determine what was tripping up the circuit.  We determined a rogue capacitor to be at fault and removed it from the circuit.

We used a spectrum analyzer to measure the frequency response of our circuit (see details in last elog).  We are currently making a Simulink block diagram so we can check the stability of our setup, but are temporarily set back because our plotted calculation of the transfer function clearly doesn't match the measured one.

ELOG V3.1.3-